About how far does the tip of a 1.8m pendulum rise as it swing 24* (degree) to one side?
Answer by Alan3354(69443)

About how far does the tip of a 1.8m pendulum rise as it swing 24* (degree) to one side?
--------------
The arc is 1.8*2*pi*(24/360) =~ 0.754 meters
The distance from the center line = 1.8*tan(24) =~ 0.732 meters
The vertical distance from the bottom center = 1.8 - 1.8*cos(24) =~ 0.1556 meters
---------------------
Pendulum swing is usually limited to ~8 degrees each side of center.